Woodcarving Co. incurred the following costs during May:

Conversion costs $ 460,000
Prime costs $ 390,000
Manufacturing overhead $ 315,000

What was the amount of direct materials and direct labor used in May?

Direct materials Direct labor
A. $ 100,000 $ 295,000
B. $ 105,000 $ 215,000
C. $ 245,000 $ 145,000
D. $ 70,000 $ 75,000
Answer #1

Prime cost=Direct material+Direct labor

Conversion cost=Direct labor+Manufacturing overhead

460,000=Direct labor+315,000

Direct labor=(460,000-315,000)=$145,000

390,000=Direct material+145,000

Direct material=(390,000-145,000)=$245,000

Hence the correct option is C.